What is the value of spin only magnetic moment of fe2+ ion?

Grade:12

1 Answers

Mehersha
6 years ago
Fe:4s^2 3d^6 and Fe^2+:3d^6. magnetic moment=root of n(n+2)=root of 4(4+2)=4.89[n=4 since it has 4 unpaired electrons]